OpenStudy (anonymous):

what is the mass of an object that is accelerated at 30 m/s^2 by a force of 150 N ?

OpenStudy (anonymous):

use formula f= ma where f=force=150N a=acceleration=30m/s^2 m=f/a=150/30=5 kg

OpenStudy (anonymous):

thank you